# LIST: Learning to Index Spatio-Textual Data for Embedding based Spatial Keyword Queries

### v2

@@ 6 Conclusions and Future Work
In this paper, we propose a novel retriever LIST, which comprises a new relevance model and a novel index. We conduct extensive experiments to show the effectiveness and efficiency of LIST over the state-of-the-art baselines. This work opens up a promising research direction of designing novel ANNS indexes for accelerating the search for embedding based spatial keyword queries. One interesting future direction is integrating our index with product quantization techniques to further expedite the search process. Another potential direction is to extend our proposed index to vector databases for dense vectors without spatial information.

### v3

@@ 6 Conclusions and Future Work
In this paper, we propose a novel index LIST for embedding based spatial keyword queries. We conduct extensive experiments to show the effectiveness and efficiency of LIST over the state-of-the-art index baselines. This work opens up a promising research direction of designing novel ANNS indexes for accelerating the search for embedding based spatial keyword queries. One interesting future direction is integrating our index with product quantization techniques to further expedite the search process. Another potential direction is to extend our proposed index to vector databases for dense vectors without spatial information.